# “Sharing, Not Showing Off”: How BeReal Approaches Authentic Self-Presentation on Social Media Through Its Design

### v1

@@ 2.3. Challenges with Authentic Self-Presentation Online
While authentic self-presentation is highly valued, particularly among teenagers, it can be particularly challenging to achieve on social media. One characteristic of computer-mediated communication (CMC) that makes it more conducive to self-presentation than face-to-face (F2F) interactions is its asynchronicity (Schlosser, 2020). In contrast to the often messy, unscripted, and spontaneous nature of F2F communication, social media allows for more controlled exchanges due to the ability to reflect, edit, and revise messages (turkle2016reclaiming). Though self-presentation can also be spontaneous (Schlenker and Wowra, 2003), self-presentation is a complex process, reflecting a dynamic interplay between the self and the audience within a specific social context (Schlenker, 2003). It involves not only expressing oneself but also responding to situational pressures and conforming to the identity expectations of significant others (Schlenker et al., 1996). It embodies aspects of an individual’s self-concept, personality, social roles, and perceptions of audience preferences. Consequently, people often leverage the asynchronicity of CMC to craft their self-presentation online (vazire2004perceptions) carefully.
@@
Online communication also differs from F2F interactions because it typically involves broadcasting messages to many people simultaneously rather than engaging in one-on-one exchanges (Schlosser, 2020). This broadcasting nature of social media interacts with adolescent characteristics in ways that significantly complicate self-presentation. Adolescents typically experience a heightened sense of self-consciousness relative to adults or younger children, leading to feelings of being under constant scrutiny, a phenomenon referred to as the imaginary audience (Elkind, 1967). This feeling is amplified in online environments where audiences are typically larger and less transparent than in offline scenarios (Zheng et al., 2019). Moreover, adolescents’ imagined audience on social media, defined as “the mental conceptualization of the people with whom [they] are communicating“ (Litt, 2012), encompasses individuals from diverse contexts such as family and school, mirroring the complex layers of audiences they engage with on these platforms. This sense of context collapse (Boyd, 2008b) prompts adolescents to employ self-presentation strategies in online environments to manage the tensions that arise (Marwick and Boyd, 2011; Boyd, 2008a; Tao and Ellison, 2023; Vitak, 2012). These strategies often involve disengaging with authentic self-presentation and instead sharing the “lowest common denominator” (Hogan, 2010): content that is universally deemed appropriate by all groups within the imagined audience (Hogan, 2010; Ullman, 1987). Consequently, relationship-building processes are also affected, as the breadth, depth, and reciprocity of authentic self-disclosure become restricted, and users disclose only superficial aspects of themselves (Hogan, 2010) or disengage entirely from sharing (Hargittai and Walejko, 2008).
@@
The third distinguishing factor of many social media sites from F2F communication is the publicly visible and quantifiable nature of audience feedback, such as ‘likes’ and comments. Adolescents’ heightened self-consciousness often drives them to present themselves favorably online, striving to appear interesting, well-liked, and attractive to maintain their social standing (Yau and Reich, 2019; Haimson et al., 2021; Hamilton et al., 2022). They strive to expand their friend networks (since network size can suggest social embeddedness (Yau and Reich, 2019)), understand the social norms of the platforms (Naegele and Goffman, 1956), and create favorable profiles (often with assistance from close friends) to achieve peer acceptance (Yau and Reich, 2019). This leads to a tension between authenticity and self-presentation online known as the online authenticity paradox (Haimson et al., 2021): people aim for authenticity in their personal life and online presentation, but doing so necessitates sharing negative experiences. Sharing these experiences requires vulnerability, which can either be risky or beneficial (Barta and Andalibi, 2021). The opportunity to present oneself on social media opens up new avenues for identity development, including experimenting with self-presentation and expression (Weinstein and Davis, 2017). However, online self-disclosure often involves risks and harms (Barta and Andalibi, 2021), and adolescents, particularly those who are sensitive to rejection and self-presentation (Yau and Reich, 2019), justifiably feel more cautious about self-disclosure (Ryan and Ryan, 2019). Explicitly delivering effort investment in communication (Kelly et al., 2017) may be helpful, but interactions on current social media revolve around “Likes” which often fail to provide the necessary support to make users feel safe in those interactions (Prinstein, 2023).
@@
Collectively, this line of work suggests a need for online spaces that support adolescent peer interactions in ways that are developmentally sensitive, i.e., providing support for authentic self-disclosure that leads to relationship development. Yet, research suggests that, currently, social media platforms instead prey upon the fears that adolescents experience, such as those stemming from their heightened sense of self-consciousness (Bukowski et al., 1998). For example, photo sharing, a prevalent form of communication on social media platforms, can exacerbate self-presentation concerns (Yau and Reich, 2019; Zhao et al., 2008). The perception of attractiveness is often central to adolescents’ self-concept and self-esteem, particularly among girls. These social media sites, with their visually oriented nature, the capability to carefully curate and manipulate visual images, and the existence of a peer audience that provides reinforcing feedback (Choukas-Bradley et al., 2020), often promote a never-ending cycle of upward social comparison. This comparison is made with meticulously selected and edited photos of peers, celebrities, and “influencers” (Fardouly et al., 2017), leading to unhealthy social comparisons and an amplified emphasis on self-presentation. Furthermore, social media platforms frequently place a higher emphasis on metrics such as time on site and return visits than on the quality of relationships. This approach encourages their users to display carefully curated content for approval and partake in continuous social comparisons.
@@ 2.4. Online Authenticity and Technology Design
While certain affordances (Evans et al., 2017; Treem and Leonardi, 2013; DeVito et al., 2017) of online communication—like asynchronicity, editability, visualness, and publicness—encourage manicured self-presentation, other affordances can promote authentic self-disclosure. Platforms and features such as Snapchat, Instagram Story, and BeReal use ephemeral content, which despite potential context loss (Cavalcanti et al., 2017), grants users greater control over their narrative and privacy (DeVito et al., 2017). This ephemerality can reduce self-consciousness, encourages casual and authentic exchanges (Bayer et al., 2016; Choi et al., 2020; Kreling et al., 2021; Xu et al., 2016; Chiu and Yuan, 2021), and enables deeper interactions, playful interactions, and private engagements (Phua et al., 2017; Choi et al., 2020; Chiu and Yuan, 2021; Xu et al., 2016).
@@
Besides ephemerality, the perceived degree of audience control and the (smaller) size of the network also support authenticity. Users often set aside online spaces for closer relationships, fostering authentic interactions within intimate friend circles (Vaterlaus et al., 2016; Piwek and Joinson, 2016; Kreling et al., 2021; Chiu and Yuan, 2021; Taber and Whittaker, 2020; Huang and Vitak, 2022). Finsta users, for example, reconfigure Instagram to establish smaller, more intimate environments, enabling the sharing of unfiltered and emotionally expressive content (Xiao et al., 2020). While this active audience management can be demanding, it effectively promotes authenticity on social media (Xiao et al., 2020). Furthermore, other affordances such as pseudonymity that increase the perceived level of trust in a network provider and its users—commonly found on platforms like Tumblr—have been shown to contribute to online authenticity, particularly allowing for marginalized individuals seeking a space for authentic self-disclosure (DeVito et al., 2018; Davis, 2023; Barta and Andalibi, 2021; Schaadhardt et al., 2023).
@@
Previous studies in Human-Computer Interaction (HCI) and Computer-Supported Cooperative Work (CSCW) have also sought to explore new ways of supporting relationship-building and authentic self-disclosure online. For instance, an analysis by Stepanova et al. (Stepanova et al., 2022) investigated the design of 50 artifacts aimed at promoting sincere feelings of connection in technologically driven systems. Their research identified nine diverse design strategies that enable genuine connection, with affective self-disclosure being the most pertinent to our study. Numerous systems employing this approach have relied on biofeedback sharing (e.g., signaling physiological activity such as heart rate and its associated emotion to a partner) in dyadic communication settings (Qin et al., 2021; Mentis et al., 2014; Liu et al., 2021, 2019; Gervais et al., 2016; Frey et al., 2018). Alternative strategies cater to broader connections beyond peer relationships, such as linking with the unity of humanity (Howell et al., 2019), engaging with strangers (Mitchell and Boer, 2017), or maintaining distant relationships (Robinson et al., 2020). In addition, a substantial body of research exists that examines how technology can facilitate self-disclosure and relationship enhancement through the use of situational conversation prompts. For instance, Zheng et al. (Zheng et al., 2021) developed PocketBot, a chatbot that nudges users to resume, enliven, and deepen conversations. However, research in this area has primarily focused on using conversational prompts to support relationships with strangers (Hu, 2022).
@@
Finally, several studies have explored BeReal from authenticity and/or relationship angles (Vanhoffelen et al., 2023; Tirocchi, 2023; Bulchand-Gidumal, 2024; Snyder, 2024; Maddox, 2023). First, Tirocchi (Tirocchi, 2023) discusses how BeReal’s authenticity aligns with the values of users aged between 20 and 23, showing that young people’s media choices, including BeReal, reflect a desire for authenticity amidst a landscape of shifting social norms and expectations. Maddox (Maddox, 2023) highlights how BeReal diverges from performance-based authenticity by emphasizing the promotion of “authenticity-as-realness” through spontaneity. The authors note how the app’s panoptic aspect and self-monitoring are what enable users to put realness before performance, a culture not seen previously on other social media platforms. Taylor (Taylor, 2023) analyzed BeReal’s designs and its creation of a “live networked public,” suggesting that the platform’s unique affordances foster what is referred to as “sporadic” authenticity that is significant for its branding against other social media apps. On the other hand, Snyder (Snyder, 2024) critiqued BeReal’s claim to authenticity by examining its user experience and rules that enforce an ‘always-on’ mentality, arguing that this approach increases the need for external curation instead of promoting true self-expression. Finally, Reddy and Kumar (Reddy and Kumar, 2024) explored how BeReal’s design features shaped what constitutes “authenticity,” finding that BeReal’s affordances align more with extemporaneous interaction than comprehensive self-presentation. Despite extensive research on BeReal and online authenticity, there remains a limited understanding of how BeReal’s design aligns with adolescent developmental needs, particularly regarding self-presentation strategies and relationship-building on social media. Our study fills this gap by analyzing BeReal’s approach beyond affordances, examining how its design cultivates specific cultural norms and expectations of authenticity that resonate with teens’ heightened sensitivity to peer perceptions while also exploring how these factors contribute to both engagement and potential toxicity in self-presentation practices.
@@ 3. Background
Table with descriptions of BeReal’s key features as of December, 2022. Design Description Image Daily notifications at random times Once at any moment of each day, the app sends out a notification with the message, “Time to BeReal. 2 min left to capture a BeReal and see what your friends are up to!” Figure 1a Two-minute posting window Clicking on the daily notification initiates a two-minute countdown or an upward counting of how late the post is, shown above the camera Figure 1b & c Retake counter Users can view the number of times any photo, including a friend’s photo, was retaken before it was posted Figure 1d Ephemerality BeReal posts, comments, and reactions remain visible to other users until the subsequent notification is sent on the following day, indicating that another post is due, while the posters themselves retain continuous access. Front and back photoshoot All BeReal posts require that users capture images from both the front and rear camera almost at the same time Figure 1e & f No filter The in-the-moment pictures taken on BeReal cannot be modified, and no filters are available. Users are not able to choose past images from photo albums either. Photo feedback After the first photo is taken with the front(back) camera and while users wait for the second photo to be taken with the back(front) camera, the platform provides feedback on the initial image taken. The feedback exerts subtle pressure encouraging users to show their face and smile, for example, “Don’t move and smiiiile,” and “Wait, wait, wait, now smile.” Figure 1g Realmojis Realmojis enable users to respond to posts with a photo of their own instead of a standard “like” or “heart.” There are six options: thumbs up, happy face, surprised face, heart eyes, crying while laughing, and a lightning bolt. The first five, or “pre-made” (P18) Realmojis, utilize stored photos that users have taken at some point to represent each of these five emotions. The lightning bolt is an “Instant Realmoji” that requires a new photo each time. Figure 1h & i Reciprocal posting Users on BeReal can only view other people’s posts if they have contributed their content for the day Figure 1j Chronological feed Unlike most mainstream social media platforms, BeReal’s feeds are strictly chronological. Consequently, users see posts from all their BeReal friends, regardless of how frequently they interact or their relationship strength. Friend count hidden BeReal does not highlight the number of friends a user has. Users cannot see the friend count of others. Concerning their own friend count, they will see how many friends they have until they surpass 50 friends, at which point their friend count will be displayed as “50+.” Friction for adding new friends The platform explicitly promotes the intentional addition of friends with statements such as “On BeReal connect with your close friends” or “Add your real friends on BeReal” beneath the friend request button. Unlike other platforms, which make numerous friend suggestions, BeReal only syncs local phone contacts and does not allow user searches. Figure 1k
@@
BeReal incorporates several design elements that funnel users towards creating a space for sharing unfiltered photos with a closer circle of friends. For instance, BeReal users receive daily notifications that encourage frequent, low-stakes posting. These notifications are sent at random times each day, as determined by the system, prompting users to post photos from both their front and back cameras within a two-minute window. Although users can postpone their response to the notification and post at a later time, such posts are tagged as “late.” Notifications about these “late” posts are sent to other users, and the delay in posting is highlighted above the post, e.g., “14 hr late.” The number of retakes is also visible to friends. While taking the photo, users are nudged to show their faces with prompts like “Your friends will definitely prefer to see your face!” and “Umm, anybody here?” Users can react to posts with “Realmojis,” emojis created from photos of their own facial expressions. Users are unable to view others’ posts until they have shared their own BeReal, and posts in feeds are displayed in chronological order. BeReal does not feature like counts or friend/follower counts. While users can see their own number of friends, any number exceeding 50 is simply displayed as “50+.” The platform also explicitly promotes adding “real friends” on BeReal. A detailed overview of these features, as of December 2022, when the study interviews were conducted, is presented in Table 1.
4. Methods

### v2

@@ 2.3. Challenges with Authentic Self-Presentation Online
While authentic self-presentation is highly valued, particularly among teenagers, it can be particularly challenging to achieve on social media. One characteristic of computer-mediated communication (CMC) that makes it more conducive to self-presentation than face-to-face (F2F) interactions is its asynchronicity (Schlosser, 2020). In contrast to the often messy, unscripted, and spontaneous nature of F2F communication, social media allows for more controlled exchanges due to the ability to reflect, edit, and revise messages (Turkle, 2016). Though self-presentation can also be spontaneous (Schlenker and Wowra, 2003), self-presentation is often a complex process, reflecting a dynamic interplay between the self and the audience within a specific social context (Schlenker, 2003). It involves not only expressing oneself but also responding to situational pressures and conforming to the identity expectations of others (Schlenker et al., 1996). It embodies aspects of an individual’s self-concept, personality, social roles, and perceptions of audience preferences. Consequently, people often leverage the asynchronicity of CMC to craft their self-presentation online (Vazire and Gosling, 2004) carefully.
@@
Online communication also differs from F2F interactions because it typically involves broadcasting messages to many people simultaneously rather than engaging in one-on-one exchanges (Schlosser, 2020). This broadcasting nature of social media interacts with adolescent characteristics in ways that significantly complicate self-presentation. Adolescents typically experience a heightened sense of self-consciousness relative to adults or younger children, leading to feelings of being under constant scrutiny, a phenomenon referred to as the imaginary audience (Elkind, 1967). This feeling is amplified in online environments where audiences are typically larger and less transparent than in offline scenarios (Zheng et al., 2019). Moreover, adolescents’ imagined audience on social media, defined as “the mental conceptualization of the people with whom [they] are communicating“ (Litt, 2012), encompasses individuals from diverse contexts such as family and school, mirroring the complex layers of audiences they engage with on these platforms. This sense of context collapse (Boyd, 2008b) prompts adolescents to employ self-presentation strategies in online environments to manage the tensions that arise (Marwick and Boyd, 2011; Boyd, 2008a; Tao and Ellison, 2023; Vitak, 2012). These strategies often involve sharing the “lowest common denominator” (Hogan, 2010): content that is universally deemed appropriate by all groups within the imagined audience (Hogan, 2010; Ullman, 1987). Consequently, relationship-building processes are also affected, as the breadth, depth, and reciprocity of authentic self-disclosure become restricted, and users disclose only superficial aspects of themselves (Hogan, 2010) or disengage entirely from sharing (Hargittai and Walejko, 2008).
@@
The third distinguishing factor of many social media sites from F2F communication is the publicly visible and quantifiable nature of audience feedback, such as “Likes” and comments. The opportunity to present oneself on social media opens up new avenues for identity development, including experimenting with self-presentation and expression (Weinstein and Davis, 2017). However, online self-disclosure often involves risks and harms (Barta and Andalibi, 2021; Schlosser, 2020), and adolescents, particularly those who are sensitive to rejection and self-presentation (Yau and Reich, 2019), justifiably feel more cautious about self-disclosure (Ryan and Ryan, 2019). The heightened self-consciousness often drives adolescents to present themselves favorably online to maintain their social standing (Yau and Reich, 2019; Haimson et al., 2021; Hamilton et al., 2022). This leads to a tension between authenticity and self-presentation online known as the “online authenticity paradox” (Haimson et al., 2021): people aim for authenticity in their personal life and online presentation, but doing so necessitates sharing negative experiences. Sharing these experiences requires vulnerability, which can either be risky or beneficial (Barta and Andalibi, 2021). Explicitly investing effort in communication (Kelly et al., 2017) may be helpful, but interactions on current social media that revolve around “Likes” often fail to provide the necessary support to make users feel safe in those interactions (Canady, 2023).
@@
Collectively, this line of work suggests a need for online spaces that support adolescent peer interactions in ways that are developmentally sensitive, i.e., providing support for authentic self-disclosure that leads to relationship development. Yet, research suggests that, currently, social media platforms instead prey upon the fears that adolescents experience, such as those stemming from their heightened sense of self-consciousness (Bukowski et al., 1998). For example, photo sharing, a prevalent form of communication on social media platforms, can exacerbate self-presentation concerns (Yau and Reich, 2019; Zhao et al., 2008). The perception of attractiveness is often central to adolescents’ self-concept and self-esteem, particularly among girls. These social media sites, with their visually oriented nature, the capability to carefully curate and manipulate visual images, and the existence of a peer audience that provides reinforcing feedback (Choukas-Bradley et al., 2020), often promote social comparison. This comparison is made with meticulously selected and edited photos of peers, celebrities, and “influencers” (Fardouly et al., 2017), leading to unhealthy social comparisons and an amplified emphasis on self-presentation. Furthermore, social media platforms frequently place a higher emphasis on metrics such as time on site and return visits than on the quality of relationships. This approach further encourages users to display carefully curated content for approval and partake in continuous social comparisons.
@@ 2.4. Online Authenticity and Technology Design
While certain affordances (Evans et al., 2017; Treem and Leonardi, 2013; DeVito et al., 2017) of online communication—like asynchronicity, editability, visualness, and publicness—encourage manicured self-presentation, other affordances can promote authentic self-disclosure. Platforms and features such as Snapchat, Instagram Story, and BeReal support ephemeral sharing, which despite potential context loss (Cavalcanti et al., 2017), grants users greater control over their narrative and privacy (DeVito et al., 2017). Ephemeral posts reduce self-consciousness, encourage casual and authentic exchanges (Bayer et al., 2016; Choi et al., 2020; Kreling et al., 2021; Xu et al., 2016; Chiu and Yuan, 2021), and enable deeper interactions, playful interactions, and private engagements (Phua et al., 2017; Choi et al., 2020; Chiu and Yuan, 2021; Xu et al., 2016).
@@
Besides ephemerality, the perceived degree of audience control and the (smaller) size of the network also support authenticity. Users often set aside online spaces for closer relationships, fostering authentic interactions within intimate friend circles (Vaterlaus et al., 2016; Piwek and Joinson, 2016; Kreling et al., 2021; Chiu and Yuan, 2021; Taber and Whittaker, 2020; Huang and Vitak, 2022). Finsta (i.e., “Fake Instagram,” or secondary Instagram accounts) users, for example, “reconfigure” (Xiao et al., 2020) Instagram to establish smaller, more intimate environments, enabling the sharing of unfiltered and emotionally expressive content (Xiao et al., 2020). While this active audience management can be demanding, it effectively promotes authenticity on social media (Xiao et al., 2020). Furthermore, pseudonymity, which can be found on platforms such as Tumblr, can foster a greater sense of trust among users and with the platform itself. This affordance has been shown to promote online authenticity, particularly benefiting marginalized individuals who seek safe spaces for genuine self-expression and disclosure (DeVito et al., 2018; Davis, 2023; Barta and Andalibi, 2021; Schaadhardt et al., 2023).
@@
Previous studies in Human-Computer Interaction (HCI) and Computer-Supported Cooperative Work (CSCW) have also sought to explore new ways of supporting relationship-building and authentic self-disclosure online. For instance, an analysis by Stepanova et al. (Stepanova et al., 2022) investigated the design of 50 artifacts aimed at promoting “genuine” feelings of connection in technologically driven systems. Their research identified nine design strategies that enable genuine connection, with affective self-disclosure being the most pertinent to our study. Numerous systems employing this approach have leveraged biofeedback sharing (e.g., signaling physiological activity such as heart rate and its associated emotion to a partner) in dyadic communication settings (Qin et al., 2021; Mentis et al., 2014; Liu et al., 2021, 2019; Gervais et al., 2016; Frey et al., 2018). Alternative strategies were catered to broader connections beyond peer relationships, such as linking with the unity of humanity (Howell et al., 2019), engaging with strangers (Mitchell and Boer, 2017), or maintaining distant relationships (Robinson et al., 2020). In addition, a substantial body of research exists that examines how technology can facilitate self-disclosure and relationship enhancement through the use of situational conversation prompts. For instance, Zheng et al. (Zheng et al., 2021) developed PocketBot, a chatbot that nudges users to resume, enliven, and deepen conversations. However, research in this area has primarily focused on using conversational prompts to support relationships with strangers (Hu, 2022).
@@
Finally, several studies have explored BeReal from authenticity and/or relationship angles (Vanhoffelen et al., 2023; Tirocchi, 2023; Bulchand-Gidumal, 2024; Snyder, 2024; Maddox, 2023). First, Tirocchi (Tirocchi, 2023) discusses how BeReal’s authenticity aligns with the values of users aged between 20 and 23, showing that young people’s media choices, including BeReal, reflect a desire for authenticity. Maddox (Maddox, 2023) highlights how BeReal diverges from performance-based authenticity by emphasizing the promotion of “authenticity-as-realness” through spontaneity. The authors note how the app’s “panopticism” and self-monitoring are what enable users to put realness before performance, a culture not seen previously on other social media platforms. Taylor (Taylor, 2023) analyzed BeReal’s designs, finding that they encourage users to engage with “sporadic” authenticity. On the other hand, Snyder (Snyder, 2024) critiqued BeReal’s claim to authenticity by examining its user experience and rules that enforce an “always-on mentality,” arguing that this approach increases the need for external curation instead of promoting true self-expression. Finally, Reddy and Kumar (Reddy and Kumar, 2024) explored how BeReal’s design features shaped what constitutes “authenticity,” finding that BeReal’s affordances align more with “extemporaneous interaction” than “comprehensive self-presentation.” Despite extensive research on BeReal and online authenticity, there remains a limited understanding of how BeReal’s design aligns with adolescent developmental needs, particularly regarding self-presentation strategies and relationship-building on social media. Our study fills this gap by analyzing BeReal’s approach beyond affordances, examining how its design cultivates specific cultural norms and expectations of authenticity that resonate with teens’ heightened sensitivity to peer perceptions. We also explore how these factors contribute to both engagement and potential toxicity in self-presentation practices.
@@ 3. Background
Table with descriptions of BeReal’s key features as of December, 2022. Design Description Image Daily notifications at random times Once at any moment of each day, the app sends out a notification with the message, “Time to BeReal. 2 min left to capture a BeReal and see what your friends are up to!” Figure 1a Two-minute posting window Clicking on the daily notification initiates a two-minute countdown or an upward counting of how late the post is, shown above the camera Figure 1b & c Retake counter Users can view the number of times any photo, including a friend’s photo, was retaken before it was posted Figure 1d Ephemerality BeReal posts, comments, and reactions remain visible to other users until the subsequent notification is sent on the following day, indicating that another post is due, while the posters themselves retain continuous access. Front and back photoshoot All BeReal posts require that users capture images from both the front and rear camera almost at the same time Figure 1e & f No filter The in-the-moment pictures taken on BeReal cannot be modified, and no filters are available. Users are not able to choose past images from photo albums either. Photo feedback After the first photo is taken with either the front or back camera, users wait for the second photo to be taken with the opposite camera. During this interval, the platform provides feedback on the initial image. The feedback exerts subtle pressure encouraging users to show their face and smile, for example, “Don’t move and smiiiile,” and “Wait, wait, wait, now smile.” Figure 1g Realmojis Realmojis enable users to respond to posts with a photo of their own instead of a standard “like” or “heart.” There are six options: thumbs up, happy face, surprised face, heart eyes, crying while laughing, and a lightning bolt. The first five, or “pre-made” (P18) Realmojis, utilize stored photos that users have taken at some point to represent each of these five emotions. The lightning bolt is an “Instant Realmoji” that requires a new photo each time. Figure 1h & i Reciprocal posting Users on BeReal can only view other people’s posts if they have contributed their content for the day Figure 1j Chronological feed Unlike most mainstream social media platforms, BeReal’s feeds are strictly chronological. Consequently, users see posts from all their BeReal friends, regardless of how frequently they interact or their relationship strength. Friend count hidden BeReal does not highlight the number of friends a user has. Users cannot see the friend count of others. Concerning their own friend count, they will see how many friends they have until they surpass 50 friends, at which point their friend count will be displayed as “50+.” Friction for adding new friends The platform explicitly promotes the intentional addition of friends with statements such as “On BeReal connect with your close friends” or “Add your real friends on BeReal” beneath the friend request button. Unlike other platforms, which make numerous friend suggestions, BeReal only syncs local phone contacts and does not allow user searches. Figure 1k
@@
BeReal incorporates several design elements that funnel users towards creating a space for sharing unfiltered photos with a closer circle of friends. For instance, BeReal users receive daily notifications that encourage frequent, low-stakes posting. These notifications are sent at random times each day, as determined by the system, prompting users to post photos from both their front and back cameras within a two-minute window. Although users can postpone their response to the notification and post at a later time, such posts are tagged as “late.” Notifications about these “late” posts are sent to other users, and the delay in posting is highlighted above the post, e.g., “14 hr late.” The number of retakes is also visible to friends. While taking the photo, users are nudged to show their faces with prompts like “Your friends will definitely prefer to see your face!” and “Umm, anybody here?” Users can react to posts with “Realmojis,” emojis created from photos of their own facial expressions. These two-photo posts are called “BeReals,” and users are unable to view others’ posts until they have shared their own BeReal, and posts in feeds are displayed in chronological order. BeReal does not feature like counts or friend/follower counts. While users can see their own number of friends, any number exceeding 50 is simply displayed as “50+.” The platform also explicitly promotes adding “real friends” on BeReal. A detailed overview of these features, as of December 2022, when the study interviews were conducted, is presented in Table 1.
4. Method